OpenStudy (anonymous):

Please, have to submit my worksheet tomorrow, need help!!!! Factor the following fractions into two pairs of brackets: 3-p squared- 2p Please please please!!!

OpenStudy (anonymous):

\[(3-p+\sqrt{2p})(3-p-\sqrt{2p})\]

OpenStudy (anonymous):

If you meant\[(3-p)^2 -2p\]

jhonyy9 (jhonyy9):

i think 3-p^2 -2p = -(p^2 +2p -3)=-(p+3)(p-1)

OpenStudy (anonymous):

no not the WHOLE squared it's just the 3- p squared

OpenStudy (anonymous):

3- p2- 2p

OpenStudy (anonymous):

This one ugghh